# Goal-Oriented Multiple Access Connectivity for Networked Intelligent Systems

### v1

@@ Abstract
We design a self-decision goal-oriented multiple access scheme, where sensing agents observe a common event and individually decide to communicate the event’s attributes to the monitoring agents, to satisfy a certain goal. Decisions are based on the usefulness of contents, which are generated under uniform, change- and semantics-aware content acquisition, as well as statistics and contents of other agents. We obtain optimal activation probabilities and threshold criteria for decision-making under all schemes, maximizing a grade of effectiveness metric. Combined with a semantics-aware acquisition scheme, the self-decision scheme offers, on average, $29.52\%$ higher effectiveness, $25.13\%$ fewer drop-offs, and $67.21\%$ fewer transmissions.
@@ I Introduction
Goal-oriented semantic communication holds great promise for realizing resource-efficient and scalable networked intelligent systems [1, 2, 3], where sensing agents observe events and convey the events’ attributes to monitoring agents to achieve a certain goal. Therein, content acquisition, data transmission, and information usage become effective in satisfying the goal when only useful or significant content is considered in the communication lifecycle. The challenge is exacerbated when the medium is shared, and agents have to communicate over multiple access channels. Conventional medium access protocols, such as typed-based ALOHA, grant-based, and grant-free access, are primarily goal- and effectiveness-agnostic, sending all frames regardless of their goal-dependent usefulness and impact at the endpoint. Despite some prior work [4, 5, 6, 7, 8, 9, 10], the problem of goal-oriented multiple access and content provision remains mostly unexplored.
In this letter, we propose a self-decision goal-oriented multiple access scheme, in which each sensing agent decides to speak up or remain silent based on the usefulness of its generated contents and the statistics of the other agents and their generated contents for achieving a specific goal. The design involves computing the optimal activation probabilities of the agents and a decision-making criterion, which maximizes a grade of effectiveness metric that integrates the discrepancy error between actual and reconstructed events, the consumed resources, and the usefulness of perceived contents. Our results reveal that the proposed self-decision scheme combined with semantics-aware acquisition offers higher effectiveness with fewer content transmissions and channel drop-offs.
@@ II System Model
We consider $K$ intelligent sensing agents (ISAs) that observe a common event (information source) and convey their noisy observations to $M$ networked monitoring agents (NMAs) over a multiple access wireless channel prone to errors and universal frequency reuse (see Fig. 1). Communication occurs over independent time-slotted service periods. At each service period, all NMAs query for new observations from the ISAs in the downlink, and the ISAs respond to those queries in the uplink, providing content within the same service period. The evolution of the source $X_{t}$ at the $t$-th service period, with $t\in\mathbb{N}$, belongs to a finite state space. At each state, the event is described via a set of information attributes, defined by $\mathcal{A}=\{\mathcal{X}_{n}~{}|~{}n=1,2,…,N\}$. Here, $\mathcal{X}_{n}\in\{x_{i}^{(n)}~{}|~{}i=1,2,…,I_{n}\}$ is a level-$n$ attribute, which is notated as $\text{AL}_{n}$ and modeled by a discrete-time Markov chain (DTMC) with $I_{n}\geq 2$ states, as shown in Fig. 1. In the DTMC modeling the $\text{AL}_{n}$, we assume $p_{i,i}=q_{n}$, and $p_{i,i^{\prime}}=p_{n}$, $\forall i\neq i^{\prime}$, for $i,i^{\prime}=1,2,...,I_{n}$. Therefore, we have $q_{n}+(I_{n}\!-\!1)p_{n}=1$.
In such model, at the $t$-th service period, the NMAs query for a group of attributes $\widehat{\mathcal{A}}_{t}\subseteq\mathcal{A}$ represented by the sequence $\langle\widehat{\mathcal{A}}_{t}\rangle$, where $|\widehat{\mathcal{A}}_{t}|\leq N$. Either queried attribute is independently communicated in one slot $j$, $j=1,2,…,|\widehat{\mathcal{A}}_{t}|$, with the same order as in $\langle\widehat{\mathcal{A}}_{t}\rangle$. The levels of queried attributes are determined in line with the goal to which the NMAs subscribe. The $k$-th ISA, where $k=1,2,…,K$, can observe only a subset of attributes $\mathcal{A}_{k}\subseteq\mathcal{A}$ and generate content from its noisy observations via an encoder, yielding a codebook shared with the other agents. Moreover, $\mathcal{K}_{n}=\langle k:\mathcal{X}_{n}\in\mathcal{A}_{k}\rangle$ is the sequence of ISAs that can observe the $n$-th attribute. To estimate the event, the NMAs need to jointly decode all queried attributes and construct $\widehat{X}_{t}$ at the $t$-th service period. An attribute is correctly perceived if at least $\lceil\frac{M+1}{2}\rceil$ NMAs successfully decode it.
II-A Goal-Oriented Content Provision
At the time slot order of the $n$-th queried attribute, the $k$-th ISA, $\forall k\in\mathcal{K}_{n}$, generates content from its observation with probability $\beta_{k,n}$ under the following acquisition schemes [11]:
@@ II-A Goal-Oriented Content Provision
    Uniform: Content is generated periodically at each time slot, regardless of the prior state of the source and the latest perceived content at the NMAs.
@@
    Change-aware: Content for an attribute is generated if the state of that attribute changes at a time slot, regardless of the latest perceived content.
@@
    Semantics-aware: Content generation for an attribute is triggered once there is a discrepancy between the current actual state of the attribute and that of the latest decoded one of the same level at the NMAs.
Then, the ISA follows a self-decision policy, thanks to which it individually decides whether to transmit that content relevant to the $n$-th attribute with probability $\alpha_{k,n}$. Content is assigned a meta-value according to its usefulness in satisfying the subscribed goal [12]. We consider $v_{k,i}^{(n)}$, $\forall k,i,n$, the meta-value for the realization $x_{i}^{(n)}$ at the $k$-th ISA, and assume it is assigned based on a knowledge fusion model, as in [12, Section III-A]. In this work, the self-decision scheme takes the form of a threshold-based decision policy. Thus, the $k$-th ISA speaks up at the $j$-th slot if $v_{k,j}^{(n)}>v_{{\rm th},k}^{(n)}$; otherwise, it remains silent. The threshold $v_{{\rm th},k}^{(n)}$, shows a variable decision-making criterion for the $k$-th ISA observing the $n$-th attribute.
@@ II-B Transmission Success Probability
As observations are noisy, contents generated from the same event via different ISAs can be different, which leads to signal contamination over the wireless network. We consider $h_{k,m}=\bar{h}_{k,m}d_{k,m}^{-\frac{a}{2}}$ the channel between the $k$-th ISA and the $m$-th NMA, represented by an independent and identically distributed (i.i.d.) random variable (r.v.) where $\bar{h}_{k,m}$ is the small-scale fading with its power having a cumulative distribution function (CDF) of $F_{h^{2}}(|\bar{h}_{k,m}|^{2})=1-\operatorname{exp}(-|\bar{h}_{k,m}|^{2})$, $d_{k,m}$ denoting the i.i.d. distance r.v.s, and $72\leq a<$ 7 being the path-loss exponent. Let us consider $\gamma_{\rm th}$ the signal-to-noise ratio (SNR) sensitivity. The probability of successful transmission for content sent from the $k$-th ISA and received at all NMAs is $\gamma_{k,n}^{(\mathcal{C},\mathcal{I})}=\mathrm{Pr}\big{(}\operatorname{SNR}_{k,n}^{(\mathcal{C},\mathcal{I})}>\gamma_{\rm th}\big{)}$. In this regard, the channel drop-off occurs when $\operatorname{SNR}_{k,n}^{(\mathcal{C},\mathcal{I})}\leq\gamma_{\rm th}$, with the probability of $1-\gamma_{k,n}^{(\mathcal{C},\mathcal{I})}$. Here, $\mathcal{C}$ shows the set of collaborating ISAs, and $\mathcal{I}$ is that of interfering ones. After some algebraic manipulations, we reach
$\gamma_{k,n}^{(\mathcal{C},\mathcal{I})}$ $=\sum_{m=1}^{M}\!\Bigg{[}\sum_{k^{\prime}\in\mathcal{C}_{k}}\sum_{k^{\prime\prime}\in\mathcal{I}}\dfrac{\prod_{k^{\prime}_{0}\in\mathcal{C}_{k}}\Lambda_{k^{\prime}_{0},m}^{(n)}\prod_{k^{\prime\prime}_{0}\in\mathcal{I}}\Omega_{k^{\prime\prime}_{0},m}^{(n)}}{\!\left(\Lambda_{k^{\prime},m}^{(n)}\!+\!\Omega_{k^{\prime\prime},m}^{(n)}\right)\!\Lambda_{k^{\prime},m}^{(n)}\Psi_{k,k^{\prime},k^{\prime\prime},m}^{(n)}}$
@@
$\Psi_{k,k^{\prime},k^{\prime\prime},m}^{(n)}=\!\!\prod_{i\in\mathcal{C}_{k}\setminus\{k^{\prime}\}}\!\!\!\left(\Lambda_{i,m}^{(n)}\!-\!\Lambda_{k^{\prime},m}^{(n)}\right)\!\times\!\!\prod_{j\in\mathcal{I}\setminus\{k^{\prime\prime}\}}\!\!\!\left(\Omega_{j,m}^{(n)}\!-\!\Omega_{k^{\prime\prime},m}^{(n)}\right)\!,$
@@ III Performance Metrics
We introduce a new metric, named grade of effectiveness (GoE), which measures the effectiveness of the content communicated over the multiple access network to satisfy the subscribed goal. This metric is defined as $\mathcal{G}=\varphi(g(\mathcal{I}))$, where $g:\mathbb{R}^{c}\to\mathbb{R}^{c^{\prime}},c\geq c^{\prime}$, is a (nonlinear) function of $c\in\mathbb{Z}^{+}$ content effectiveness features $\mathcal{I}\in\mathbb{R}^{c}$, and $\varphi:\mathbb{R}^{c^{\prime}}\to\mathbb{R}$ is a context-dependent, cost-aware function [2].
In this letter, we consider the following three effectiveness features: effective discrepancy error (EDE), effective resource consumption (ERC), and effective utility of content (EUoC).
@@ III-A Effective Discrepancy Error (EDE)
A content discrepancy error $E_{t}$ occurs at the $t$-th service period if the actual event $X_{t}$ is not equal to the reconstructed $\widehat{X}_{t}$, i.e., $E_{t}=\mathbbm{1}\{X_{t}\neq\widehat{X}_{t}\}$. The EDE depicts the feature that the effect of losing information is scaled by its usefulness in satisfying the goal. Given this, we model the EDE as follows
${\rm EDE}(E_{t},\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})=\mathcal{F}_{1}\sum_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}P_{{\rm e},n}$ (2)
where $\text{where~{}}\mathcal{F}_{1}$ $=\sum_{j=1}^{{\mid\widehat{\mathcal{A}}_{t}\mid}}\left[\,\sum_{k\in\mathcal{K}_{n}}\alpha_{k,n}v_{k,j}^{(n)}\right]_{n\leftarrow\mathcal{A}_{j}},$
$\boldsymbol{\alpha}=[\alpha_{k,n}]_{K\times N}$, $\mathbf{v}=[v_{k,j}^{(n)}]_{K\times\mid\widehat{\mathcal{A}}_{t}\mid\times N}$, $\mathcal{A}_{j}\subseteq\widehat{\mathcal{A}}_{t}$ shows the attribute queried at the $j$-th time slot, and the notation $n\leftarrow\mathcal{A}_{j}$ implies that $\mathcal{A}_{j}$ is the $n$-th attribute. We also have
$P_{{\rm e},n}=\underset{T\to\infty}{\lim}\frac{1}{T}\sum_{t=1}^{T}\bigg{[}\mathrm{Pr}\!\left({E}_{t,j}^{(n)}=1\right)\!\bigg{]}_{j:\mathcal{A}_{j}\in\widehat{\mathcal{A}}_{t}},$ (3)
@@
where $\mathcal{E}_{n}$ denotes the probability that the $n$-th attribute is not successfully delivered to at least $\lceil\frac{M+1}{2}\rceil$ NMAs, defined as
${\mathcal{E}}_{n}=\sum_{\ell=1}^{2^{|\mathcal{K}_{n}|}}\tilde{\mathcal{E}}_{n,\ell}$ (5)
with ~$\tilde{\mathcal{E}}_{n,\ell}$ given by
~$\tilde{\mathcal{E}}_{n,\ell}=\prod_{r^{\prime}\in\mathcal{R}^{\prime}_{\ell}}(1\!-\!\alpha_{r^{\prime},n})\prod_{r\in\mathcal{R}_{\ell}}\alpha_{r,n}$
×$~{}\times\sum_{\jmath=1}^{2^{|\mathcal{R}_{\ell}|}}\Bigg{[}\prod_{\tau^{\prime}\in\mathcal{T}^{\prime}_{\jmath}}(1\!-\!\chi_{\tau^{\prime},n})\prod_{\tau\in\mathcal{T}_{\jmath}}\chi_{\tau,n}\prod_{v\in\mathcal{V}_{\jmath}}\!\left(1\!-\!\gamma_{v,n}^{(\mathcal{T}_{\jmath},\mathcal{T}_{\jmath}^{\prime})}\right)\!\Bigg{]}$ (6)
where $\chi_{k,n}$ is the probability that the $k$-th ISA’s observation is correct. Also, $\mathcal{R}_{\ell}$, $\mathcal{R}^{\prime}_{\ell}$, $\mathcal{T}_{\jmath}$, and $\mathcal{T}^{\prime}_{\jmath}$ are the sets of ISAs that are active, inactive, active with correct observations, and active with incorrect observations, respectively. Moreover, $\mathcal{V}_{\jmath}$ shows the subset of $\lceil\frac{M}{2}\rceil$ ISAs from $\mathcal{T}_{\jmath}$ with the longest distances from all NMAs if $\mid\!\mathcal{T}_{\jmath}\!\mid\,\geq\lceil\frac{M}{2}\rceil$, $\forall\jmath$; otherwise, $\mathcal{V}_{\jmath}$ is a null set.
@@
We use the same approach as in [11, Section 3.2] to derive (4). Considering its definition, the discrepancy error $E_{t,j}^{(n)}$, $\forall t,j,n$, follows a DTMC with two possible states ${\{0,1\}}$ and transition probabilities $\pi^{(n)}_{a,b}=\mathrm{Pr}\big{(}E_{t+1,j^{\prime}}^{(n)}=b~{}|~{}E_{t,j}^{(n)}=a\big{)}$, $\forall a,b\in\{0,1\}$, where $\mathcal{A}_{j}=\mathcal{A}_{j^{\prime}}=n$. Therefore, we can obtain the steady-state error probability $P_{{\rm e},n}$, as follows

### v2

@@ Abstract
We design a self-decision goal-oriented multiple access scheme, where sensing agents observe a common event and individually decide to communicate the event’s attributes as updates to the monitoring agents, to satisfy a certain goal. Decisions are based on the usefulness of updates, generated under uniform, change- and semantics-aware acquisition, as well as statistics and updates of other agents. We obtain optimal activation probabilities and threshold criteria for decision-making under all schemes, maximizing a grade of effectiveness metric. Alongside studying the effect of different parameters on effectiveness, our simulation results show that the self-decision scheme may attain at least $92\%$ of optimal performance.
Index Terms:
Goal-oriented multiple access, semantic update acquisition, optimal activation probability, decision-making
@@ I Introduction
[…]
Goal-oriented semantic communication holds great promise for realizing resource-efficient and scalable networked intelligent systems [1], where sensing agents observe events and convey the events’ attributes to monitoring agents to achieve a certain goal. Therein, update acquisition, data transmission, and information usage become effective in satisfying the goal when only useful or significant updates are considered in the communication lifecycle. The challenge is exacerbated when the medium is shared, and agents have to communicate over multiple access channels. Conventional medium access protocols, such as typed-based ALOHA, grant-based, and grant-free access, are primarily goal- and effectiveness-agnostic, sending all frames regardless of their goal-dependent usefulness and impact at the endpoint. Despite some prior work [2, 3, 4, 5, 6, 7], the problem of goal-oriented multiple access and update provision remains mostly unexplored.
In this letter, we propose a self-decision goal-oriented multiple access scheme, in which each sensing agent decides to speak up or remain silent based on the usefulness of its generated updates and the statistics of the other agents and their generated updates for achieving a specific goal. The design involves computing the optimal activation probabilities of the agents and a decision-making criterion, which maximizes a grade of effectiveness metric that integrates the discrepancy error between actual and reconstructed events, the consumed resources, and the usefulness of perceived updates. Our results indicate that the proposed self-decision scheme is very close to the optimal solution.
@@ II System Model
We consider $K$ intelligent sensing agents (ISAs) that observe a common event (information source) and convey their noisy observations in the form of updates to $M$ networked monitoring agents (NMAs) over a multiple access wireless channel prone to errors and universal frequency reuse, c.f., Fig. 1. Communication occurs over independent time-slotted service intervals. Within each interval, all NMAs query for new observations from the ISAs in the downlink, and the ISAs respond to those queries in the uplink, providing updates within the same service interval. The evolution of the source $X_{t}$ at the $t$-th service interval, with $t\in\mathbb{N}$, belongs to a finite state space. An event at each state is described via a set of information attributes, defined by $\mathcal{A}=\{\mathcal{X}_{n}~{}|~{}n=1,2,\dots,N\}$. Here, $\mathcal{X}_{n}\in\{x_{i}^{(n)}~{}|~{}i=1,2,\dots,I_{n}\}$ is a level-$n$ attribute, which is modeled by a discrete-time Markov chain (DTMC) with $I_{n}\geq 2$ states. In the DTMC modeling the $n$-th attribute, we assume $p_{i,i}=p_{n}^{\prime}$, and $p_{i,i^{\prime}}=p_{n}$, $\forall i\neq i^{\prime}$, for $i,i^{\prime}=1,2,\dots,I_{n}$, where $p_{n}^{\prime}+(I_{n}\!-\!1)p_{n}=1$.
@@
At the $t$-th interval, the NMAs query for a group of attributes $\widehat{\mathcal{A}}_{t}\subseteq\mathcal{A}$ represented by the sequence $\langle\widehat{\mathcal{A}}_{t}\rangle$, where $|\widehat{\mathcal{A}}_{t}|\leq N$. Either queried attribute is independently communicated in one slot $j$, $j=1,2,\dots,|\widehat{\mathcal{A}}_{t}|$, with the same order as in $\langle\widehat{\mathcal{A}}_{t}\rangle$. The levels of queried attributes are determined in line with the goal to which the NMAs subscribe. The $k$-th ISA, where $k=1,2,\dots,K$, can observe only a subset of attributes $\mathcal{A}_{k}\subseteq\mathcal{A}$ and generate updates from its noisy observations via an encoder, yielding a codebook shared with the other agents. Moreover, $\mathcal{K}_{n}=\langle k:\mathcal{X}_{n}\in\mathcal{A}_{k}\rangle$ is the sequence of ISAs that can observe the $n$-th attribute. To estimate the event, the NMAs need to jointly decode all queried attributes and construct $\widehat{X}_{t}$ at the $t$-th service interval. An attribute is correctly perceived at that interval if at least $M_{t}$ NMAs successfully decode it, where $1\leq M_{t}\leq M$.
II-A Goal-Oriented Update Provision
At the time slot order of the $n$-th queried attribute, the $k$-th ISA, $\forall k\in\mathcal{K}_{n}$, generates an update from its observation with probability $\beta_{k,n}$, as obtained in Section IV-D, under the following acquisition schemes [8]:
@@ II-A Goal-Oriented Update Provision
    Uniform: An update is generated periodically at each time slot, regardless of the prior state of the source and the latest perceived update at the NMAs.
@@
    Change-aware: An update for an attribute is generated if the state of that attribute changes at a time slot, regardless of the latest perceived update.
@@
    Semantics-aware: Update generation for an attribute is triggered once there is a discrepancy between the current actual state of the attribute and that of the latest decoded one of the same level at the NMAs. In this case, the ISAs should maintain a record of the last decoded attributes.
The ISA follows a self-decision policy, thanks to which it individually decides whether to transmit that update relevant to the $n$-th attribute with probability $\alpha_{k,n}$. Each update is assigned a meta value according to its usefulness in satisfying the subscribed goal [9]. We consider $v_{k,i}^{(n)}$, $\forall k,i,n$, the meta value for the realization $x_{i}^{(n)}$ at the $k$-th ISA, and assume it is assigned based on a knowledge fusion model, as in [9, Section III-A]. In this work, the self-decision scheme takes the form of a threshold-based decision policy. Thus, the $k$-th ISA speaks up at the $j$-th slot if $v_{k,j}^{(n)}>v_{{\rm th},k}^{(n)}$; otherwise, it remains silent. The threshold $v_{{\rm th},k}^{(n)}$, shows a variable decision-making criterion for the $k$-th ISA observing the $n$-th attribute.
@@ II-B Transmission Success Probability
As observations are noisy, updates generated from the same event via different ISAs can be different, which leads to signal contamination over the wireless network. We consider $h_{k,m}=\bar{h}_{k,m}d_{k,m}^{-\frac{a}{2}}$ the channel between the $k$-th ISA and the $m$-th NMA, represented by an independent and identically distributed (i.i.d.) random variable (r.v.) where $\bar{h}_{k,m}$ is the small-scale fading with its power having a cumulative distribution function (CDF) of $F_{h^{2}}(|\bar{h}_{k,m}|^{2})=1-\operatorname{exp}(-|\bar{h}_{k,m}|^{2})$, $d_{k,m}$ denoting the i.i.d. distance r.v.s, and $72\leq a<$ 7 being the path-loss exponent. Let us consider $\gamma_{\rm th}$ the signal-to-noise ratio (SNR) sensitivity. The probability of successful transmission of the update from the $k$-th ISA and received at the $m$-th NMA is $\gamma_{k,m,n}^{(\mathcal{C},\mathcal{I})}=\mathrm{Pr}\big{(}\operatorname{SNR}_{k,m,n}^{(\mathcal{C},\mathcal{I})}>\gamma_{\rm th}\big{)}$. The channel drop-off occurs when $\operatorname{SNR}_{k,m,n}^{(\mathcal{C},\mathcal{I})}\leq\gamma_{\rm th}$, with the probability of $1-\gamma_{k,m,n}^{(\mathcal{C},\mathcal{I})}$. Here, $\mathcal{C}$ shows the set of collaborating ISAs, and $\mathcal{I}$ is that of interfering ones. After some algebraic manipulations, we reach
$\gamma_{k,m,n}^{(\mathcal{C},\mathcal{I})}$ $=\sum_{k^{\prime}\in\mathcal{C}_{k}}\sum_{k^{\prime\prime}\in\mathcal{I}}\Bigg{[}\dfrac{\prod_{k^{\prime}_{0}\in\mathcal{C}_{k}}\Lambda_{k^{\prime}_{0},m}^{(n)}\prod_{k^{\prime\prime}_{0}\in\mathcal{I}}\Omega_{k^{\prime\prime}_{0},m}^{(n)}}{\big{(}\Lambda_{k^{\prime},m}^{(n)}\!+\!\Omega_{k^{\prime\prime},m}^{(n)}\big{)}\Lambda_{k^{\prime},m}^{(n)}\Psi_{k,k^{\prime},k^{\prime\prime},m}^{(n)}}$
@@
$\Psi_{k,k^{\prime},k^{\prime\prime},m}^{(n)}$ $=\prod_{i\in\mathcal{C}_{k}\setminus\{k^{\prime}\}}\!\!\!\left(\Lambda_{i,m}^{(n)}\!-\!\Lambda_{k^{\prime},m}^{(n)}\right)\!$
×$~{}~{}~{}\times\!\!\prod_{j\in\mathcal{I}\setminus\{k^{\prime\prime}\}}\!\!\!\left(\Omega_{j,m}^{(n)}\!-\!\Omega_{k^{\prime\prime},m}^{(n)}\right)\!,$
@@ III Performance Metrics
We propose a grade of effectiveness (GoE) metric to evaluate the effectiveness of communicated updates in the form of ${\rm GoE}=\varphi\left(\mathbf{W}^{T};g_{1}(f_{1}),\dots,g_{R}(f_{R})\right)$, which has $R\in\mathbb{Z}^{+}$ effectiveness features, with $f_{r}\in\mathbb{R}_{0}^{+}$ for $r=1,2,\dots,R$ denoting the $r$-th one. Herein, $\mathbf{W}=[w_{1},\dots,w_{R}]^{T}\in\mathbb{R}^{R}$, $0\leq w_{r}\leq 1$, indicates a weight factor, where $\mathbf{W}^{T}\mathbf{1}_{R\times 1}=1$. Also, $g_{r}:\mathbb{R}_{0}^{+}\to\mathbb{R}$ is a positive non-decreasing differentiable function, and $\varphi:\mathbb{R}^{2R}\to\mathbb{R}$ is a context-dependent utility function [1]. In this letter, we focus on three effectiveness features to determine the GoE: Effective discrepancy error (EDE), effective resource consumption (ERC), and effective utility of updates (EUU).
@@ III-A Effective Discrepancy Error (EDE)
Update discrepancy error $E_{t}$ occurs at the $t$-th service interval if the actual event $X_{t}$ is not equal to the reconstructed $\widehat{X}_{t}$, i.e., $E_{t}=\mathbbm{1}\{X_{t}\neq\widehat{X}_{t}\}$. The EDE depicts the feature that the effect of losing information is scaled by its usefulness in satisfying the goal. Given this, we model the EDE as follows
${\rm EDE}(E_{t},\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})=\mathcal{F}_{1}\sum_{n:\mathcal{X}_{n}\in\widehat{\mathcal{A}}_{t}}P_{{\rm e},n}$ (2)
where $\text{where~{}}\mathcal{F}_{1}$ $=\sum_{j=1}^{{\mid\widehat{\mathcal{A}}_{t}\mid}}\left[\,\sum_{k\in\mathcal{K}_{n}}\alpha_{k,n}v_{k,j}^{(n)}\right]_{n:\mathcal{X}_{n}=\mathcal{X}_{j}},$
$\boldsymbol{\alpha}=[\alpha_{k,n}]_{K\times N}$, $\mathbf{v}=[v_{k,j}^{(n)}]_{K\times\mid\widehat{\mathcal{A}}_{t}\mid\times N}$, $\mathcal{X}_{j}\in\widehat{\mathcal{A}}_{t}$ indicates the attribute queried at the $j$-th time slot. We also have
$P_{{\rm e},n}=\underset{T\to\infty}{\lim}\frac{1}{T}\sum_{t=1}^{T}\bigg{[}\mathrm{Pr}\!\left({E}_{t,j}^{(n)}=1\right)\!\bigg{]}_{j:\mathcal{X}_{j}\in\widehat{\mathcal{A}}_{t}},$ (3)
@@
where $\mathcal{E}_{n}=\frac{1}{2^{|\mathcal{K}_{n}|}}\sum_{\ell=1}^{2^{|\mathcal{K}_{n}|}}\mathcal{E}_{n,\ell}$ denotes the probability that the $n$-th attribute is not successfully delivered to $M_{t}$ NMAs, where
$\mathcal{E}_{n,\ell}=\prod_{k_{0}\in\mathcal{R}^{\prime}_{\ell}}(1\!-\!\alpha_{k_{0},n})\prod_{k_{1}\in\bar{\mathcal{R}}_{\ell}}\alpha_{k_{1},n}$
×$~{}~{}\times\!\sum_{\tau=1}^{2^{|\mathcal{R}_{\ell}|}}\!\Bigg{[}\prod_{k_{2}\in\mathcal{T}^{\prime}_{\tau}}(1\!-\!q_{k_{2},n})\prod_{k_{3}\in\mathcal{T}_{\tau}}q_{k_{3},n}\prod_{m\in\mathcal{M}_{k_{3}}}\!\!\left(1\!-\!\gamma_{k_{3},m,n}^{(\mathcal{T}_{\tau},\mathcal{T}_{\tau}^{\prime})}\right)\!\Bigg{]}$ (5)
where $q_{k,n}$ is the probability that the $k$-th ISA’s observation is correct. Moreover, $\mathcal{R}_{\ell}$, $\mathcal{R}^{\prime}_{\ell}$, $\mathcal{T}_{\tau}$, and $\mathcal{T}^{\prime}_{\tau}$ are the sets of ISAs that are active, inactive, active with correct observations, and active with incorrect observations, respectively. Also, $\mathcal{M}_{k_{3}}$ shows the subset of $M_{t}+1$ NMAs being farthest from the $k_{3}$-th ISA.
@@
We use the same approach as in [8, Section 3.2] to derive (4). Considering its definition, the discrepancy error $E_{t,j}^{(n)}$, $\forall t,j,n$, follows a DTMC with two possible states ${\{0,1\}}$ and transition probabilities $\pi^{(n)}_{a,b}=\mathrm{Pr}\big{(}E_{t+1,j^{\prime}}^{(n)}=b~{}|~{}E_{t,j}^{(n)}=a\big{)}$, $\forall a,b\in\{0,1\}$, where $\mathcal{X}_{j}=\mathcal{X}_{j^{\prime}}=\mathcal{X}_{n}$. Therefore, we can obtain the steady-state error probability $P_{{\rm e},n}$, as follows